# Imperceptible Protection Against Style Imitation from Diffusion Models

### v1

@@ Appendix 0.C Additional Analyses and Results
Ablation study. In Table 6, we conduct additional component analysis by omitting certain components from the full Impasto framework. Specifically, we assess the impact of 1) removing the JND-based perception map, which results in initializing $\mathbf{M}$ with random masks (w/o JND), and 2) excluding the proposed perception-aware protection (PAP), thereby relying solely on the perceptual constraint bank (w/o PAP). For the w/o JND case, in contrast to Table 3, we start with a full-component model and remove only the JND part. The results indicate that discarding the JNDs leads to a decrease in protection performance while maintaining similar image quality, corroborating the findings in Table 3. When PAP is not applied, there is a slight increase in protection performance (FID: 279.3 $\rightarrow$ 281.1) with significant compromise on image quality. These observations suggest that the integration of PAP alongside a JND-based perceptual map is crucial for achieving an imperceptible style protection framework that effectively balances protection performance with image quality.
Method Protected Image Quality Protection Performance
Impasto (Full model) 0.159 0.912 20.74 279.3
w/o JND 0.158 0.912 19.30 268.1
w/o PAP 0.199 0.890 17.57 281.1
Table 7: Evaluating the impact of varying refinement Step $P$. In this analysis, we explore the effects on performance when adjusting the number of refinement steps $P$ within the instance-wise refinement process.
Method Protected Image Quality Protection Performance
$P=100$ 0.159 0.913 19.29 278.4
$P=75$ 0.159 0.913 19.06 278.2
$P=50$ 0.158 0.913 18.19 279.7
$P=25$ 0.159 0.912 20.74 279.3
$P=10$ 0.159 0.913 18.49 274.3
@@
Table 8: Quantitative comparison of protection methods w/ and w/o Impasto, both selected for their comparable protection performance. Impasto markedly elevates the protected images’ quality while maintaining comparable levels of protection efficacy.
@@
Varying refine step in IWR. In this analysis, we investigate the effect of varying the refining step $P$ during the instance-wise refinement (IWR) process (Algorithm 1 L8-10) and assess the changes in protection performance (Table 7). We observe that though $P$ is decreased from 100 to 25, both the image quality and protection capability remain relatively consistent. This performance stability can be attributed to the initial utilization of a JND-based perceptual map $\mathcal{M}$, which necessitates only minor adjustments during the IWR step. However, a notable reduction in protection performance is shown when $P$ is reduced to 10, indicating that such a small step is insufficient for achieving optimal refinement. Consequently, we set $P=25$ as the default setting, balancing both performance efficacy and the number of optimization steps.
@@
Impasto with other protection methods. In addition to PhotoGuard [46] and AdvDM [30], we extend the application of Impasto to recent protection techniques, Anti-DreamBooth [52] and Mist [29]. Anti-DreamBooth, a variant of AdvDM, incorporates DreamBooth tuning into its optimization process. Mist utilizes a hybrid approach, combining both Encoder (e.g. PhotoGuard) and UNet-based (e.g. AdvDM) protection methods (similar to Eq. 4). We adhere to the hyperparameters specified in their official implementations. Regarding Impasto, we employ the same settings used in our application to AdvDM (Sec. 0.A.3), for both Anti-DreamBooth and Mist. As shown in Table 8, Impasto notably enhances the quality of the protected images with comparable protection performances. Fig. 11 and 12 also support the efficacy of our method; Impasto significantly diminishes visual artifacts in the protected images while preserving the original methods’ style protection ability. These experiments highlight the adaptability and efficacy of Impasto across various protection frameworks.
@@
Additional qualitative results. Fig. 13 and 14 provide additional visual comparisons, illustrating the impact of applying Impasto in conjunction with PhotoGuard [46] and AdvDM [30], respectively.
@@
We also compare the robustness on countermeasure of Impasto in Fig. 15, including JPEG compression, blurring, and noise addition. Across all these countermeasures, our method demonstrates a protection capability comparable to that of the baseline. Furthermore, in Fig. 16, we present additional qualitative results demonstrating the generalization capability of Impasto. When applied to three different personalization methods—DreamBooth, LoRA, and DreamStyler—Impasto maintains performance levels akin to those observed in their respective non-Impasto implementations. This highlights the effectiveness and adaptability of Impasto across diverse personalization techniques.
@@
Figure 11: Qualitative comparison of Anti-Dreambooth [52] with and without Impasto.
@@
Figure 12: Qualitative comparison of Mist [29] with and without Impasto.
@@
Figure 13: Qualitative comparison of PhotoGuard [46] with and without Impasto.
@@
Figure 14: Qualitative comparison of AdvDM [30] with and without Impasto.
@@
Figure 16: Qualitative comparison on generalization. Impasto does not impede protection methods’ generalization abilities across diverse personalization methods; DreamBooth [44], LoRA [17], and DreamStyler [1]

### v2

@@ Appendix D Discussions and Analyses
In Table 4, we show additional analysis by omitting certain components from the full Impasto framework. Specifically, we assess the impact of 1) removing the JND-based perception map, which results in initializing $\mathbf{M}$ with random masks (w/o JND), and 2) excluding the proposed perception-aware protection (PAP), thereby relying solely on the perceptual constraint bank (w/o PAP). For the w/o JND case, in contrast to Table 2, we start with a full-component model and remove only the JND part. The results indicate that discarding the JNDs leads to a decrease in protection performance while maintaining similar image quality, corroborating the findings in Table 2. When PAP is not applied, there is a slight increase in protection performance (FID: 279.3 $\rightarrow$ 281.1) with significant compromise on image quality. These observations suggest that the integration of PAP alongside a JND-based perceptual map is crucial for achieving an imperceptible style protection framework that effectively balances protection performance with image quality.
@@
Table 8: Quantitative comparison of protection methods w/ and w/o Impasto, both selected for their comparable protection performance. Impasto markedly elevates the protected images’ quality while maintaining protection efficacy.
@@
Varying refinement interval step. In this analysis, we explore the effect of varying the refinement interval step $P$ on instance-wise refinement (IWR) and difficulty-aware protection (DAP) to evaluate changes in protection performance (Table 5). We observe that even as $P$ increases from 1 to 4 (i.e., with less frequent refinements), both image quality and protection capability remain relatively stable. This stability can be attributed to the initial use of a JND-based perceptual map $\mathcal{M}$, which requires only minor adjustments during the IWR step. However, when $P$ is increased to 10, a significant reduction in protection performance is observed, indicating that such a large interval is insufficient for optimal refinement. Therefore, we set $P=4$ as the default, balancing performance efficacy with the number of optimization steps.
Table 9: Comparison on the blur countermeasure of PhotoGuard with and without Impasto at a similar level of protected image quality.
FID ($\uparrow$) PhotoGuard + Impasto
Clean 274.4 279.3
Blur 273.5 274.6
Countermeasures. As discussed earlier, Impasto achieves similar protection performance against JPEG compression and Gaussian noise countermeasures but shows slight degradation under Gaussian blur. The reason for this phenomenon is that the blurring effect removes most high-frequencies, making weak perturbations less effective. We note that if baseline protection models are applied with weak protection strength (to match the protected images’ perceptual quality to ours), they also exhibit degraded protection performance.
To demonstrate this, we match the protected image quality of PhotoGuard to ours and compare the protection performance in both clean (no countermeasure) and blur scenarios (Table 9). In this case, which with similar imperceptibility, Impasto shows better performance under blur, implying that such degradation is an inevitable limitation of using weaker perturbations rather than a flaw in our model design. Note that the performance gap narrows under blur because Impasto applies very weak perturbations to more noticeable regions, potentially causing minor vulnerabilities. Nevertheless, in all scenarios—whether clean or under countermeasures—Impasto demonstrates an improved trade-off between protected image quality and protection performance, highlighting its practical effectiveness.
Protection performance comparison. Here, unlike Table 1, which compares the protected images’ quality under similar protection performance, we compare the protection performance while fixing the image quality. In Table 6, Impasto improves protection performance compared to the baseline in most metrics and scenarios. This was also shown in the quality-protection trade-off (Figure 7). These results indicate that attaching Impasto to the protection baselines enables them to be better protection methods.
Impasto in other domain. Although Impasto is initially proposed to prevent style imitation, its applicability can be extended beyond other domains or applications. To validate this, we conduct protection on two facial datasets, CelebA-HQ (Karras et al. 2017) and VGGFace2 (Cao et al. 2018). As demonstrated in Table 7 and Figure 12, adopting Impasto in these natural domains can also enhance the quality of protected images without compromising the protection performance of baseline models. This broad applicability highlights Impasto’s versatility and potential as a universal tool for protecting users’ copyright and preventing serious threats of deepfakes.
Appendix E Additional Results
@@
Impasto with other protection methods. In addition to PhotoGuard (Salman et al. 2023) and AdvDM (Liang et al. 2023), we extend the application of Impasto to recent protection techniques, Anti-DreamBooth (Van Le et al. 2023) and Mist (Liang and Wu 2023). Anti-DreamBooth, a variant of AdvDM, incorporates DreamBooth tuning into its optimization process. Mist utilizes a hybrid approach, combining both Encoder (e.g. PhotoGuard) and UNet-based (e.g. AdvDM) protection methods (similar to Eq. 4). We adhere to the hyperparameters specified in their official implementations. Regarding Impasto, we employ the same settings used in our application to AdvDM (Section B), for both Anti-DreamBooth and Mist. As shown in Table 8, Impasto notably enhances the quality of the protected images with comparable protection performances. Figure 13 and 14 also support the efficacy of our method; Impasto significantly diminishes visual artifacts in the protected images while preserving the original methods’ style protection ability. These experiments highlight the adaptability and efficacy of Impasto across various protection frameworks.
@@
Qualitative results. Figure 15 and 16 provide additional visual comparisons, illustrating the impact of applying Impasto in conjunction with PhotoGuard (Salman et al. 2023) and AdvDM (Liang et al. 2023), respectively.
@@
Robustness and generalization. We compare the robustness on countermeasure of Impasto in Figure 17, including JPEG compression, blurring, and noise addition. Across all these countermeasures, Impasto demonstrates a protection capability comparable to that of the baseline. In Figure 18, we present additional results demonstrating the generalization capability of Impasto. When applied to three different personalization methods—DreamBooth, LoRA, and DreamStyler—Impasto maintains performance levels akin to those observed in their respective non-Impasto implementations. This highlights the effectiveness and adaptability of Impasto across diverse personalization techniques.
@@
Figure 13: Qualitative comparison of Anti-Dreambooth (Van Le et al. 2023) with and without Impasto. Best viewed in zoom.
@@
Figure 14: Qualitative comparison of Mist (Liang and Wu 2023) with and without Impasto. Best viewed in zoom.
@@
Figure 15: Qualitative comparison of PhotoGuard (Salman et al. 2023) with and without Impasto. Best viewed in zoom.
@@
Figure 16: Qualitative comparison of AdvDM (Liang et al. 2023) with and without Impasto. Best viewed in zoom.
@@
Figure 18: Qualitative comparison on generalization. Impasto does not impede baseline’s generalization abilities across diverse personalization methods; DreamBooth (Ruiz et al. 2023), LoRA (Hu et al. 2021), and DreamStyler (Ahn et al. 2023)